### v1

# Charged excitations made neutral: N-centered ensemble density functional theory of Fukui functions

Abstract
We introduce an in-principle exact working equation to compute Fukui functions within $N$-centered ensemble DFT. It avoids the kernel derivative discontinuity problem of DFT for fractional number of electrons, whose contribution is recovered through weight-derivatives of the ensemble density functional potential. We explore practical strategies to compute its contribution by recycling ground-state density-functional approximations dressed with a weight-dependent scaling function. We also show interpolating between known limits of the ensemble functional and learning from uniform density profiles is a very effective strategy.
Introduction – No matter how much we succeed in making accurate computer simulations of chemical reactions routine, there will always be a need for chemical descriptors of reactivity. While they cannot systematically replace simulation nor experiment, their appeal is to bring ”primitive patterns of understanding” [^1] Attributed to C.A. Coulson by R. McWeeny in the preface of McWeeny R (1979) Coulson’s valence. Oxford University Press, Oxford reducing the over-abundant information generated in Ab Initio resolution of so many processes to key principles we can grasp. In a time where accurate molecular simulations are not so routinely feasible, their guidance is as valuable at the prediction stage than for ensuing rationalization.
As chemistry and condensed-matter physics is first and foremost dictated by the response of electrons to perturbation (from other electrons and nuclei, light, applied electric field and many more), they are the focus of properties descriptors. The advent of Density Functional Theory (DFT), an in-principle exact framework in which every property of the system follows from the electronic density, brought a solid foundation from which sharp definitions of insightful concepts – such as electronegativityParr (1980) and hardnessParr and Pearson (1983) – emerged. Further analysis of increasingly local responses of the electronic energy to changes in the external potential and number of electrons bloomed into the field of Conceptual Density Functional TheoryGeerlings et al. (2003); Geerlings and De Proft (2008) (CDFT). In turn, their computation from corresponding functional derivatives of the (to-be-approximated) universal functional proved instrumental in understanding what information practical Density Functional Approximations (DFA) should incorporate: By focusing efforts on reproducing more relevant properties, DFA design can (hopefully) find shortcuts on the meandering path towards the exact functionalGould (2017). This is a contribution of the present work.
We will focus more specifically on the Fukui function, introduced by Parr and YangParr and Yang (1984) as a tool to understand the relative reactivity of different sites in a moleculeAyers and Levy (2000). It is defined as the derivative of the electronic density with respect to the electron number, a somewhat puzzling concept for isolated moleculesNguyen-Dang et al. (1982) which must have an integer number of electrons, but very natural for atoms or fragments within a bigger molecular arrangement resulting in partial charges on each subunitBader (1990). Based on the reasoning that an approaching reagent will prefer the site most favorable to addition/removal of an electron for a nucleophilic/electrophilic attack, the corresponding Fukui function will reach a higher value the more reactive a site isAyers and Levy (2000). It offers a (sometimes criticalBartolotti and Ayers (2005)) improvement over frontier molecular orbital theoryFukui et al. (1952, 1954); Fukui (1982) by inclusion of orbital relaxation effects.
Interpretation and computation of derivatives with respect to the number of electrons analytically, i.e. connecting them to the $N$-electron Kohn Sham system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and NN-1 electron systemsYang et al. (2012). To that end, extension of DFT to fractional electron numbers by Perdew, Parr, Levy and Balduz (PPLB) showed the generalized energy functional possesses derivative discontinuities when crossing an integer number of electronsPerdew et al. (1982), a behavior which in turn affects properties obtained through differentiating itPerdew et al. (1982); Mori-Sánchez et al. (2008); Cohen et al. (2008a, b); Stein et al. (2010); Zheng et al. (2011); Cohen et al. (2011); Perdew et al. (2017). Accurate predictions depend on inclusion of discontinuous shifts of the Hxc fields at integer electron number thresholds, which practical (local and semi-local) DFAs lack. The most well-known case is the missed contribution of Hxc potential derivative discontinuities resulting in systematic underestimation of fundamental gaps in solidsPerdew et al. (2017), also relevant for molecular dissociation and charge transferHodgson et al. (2017); Kraisler et al. (2021). It is less widely appreciated that similar discontinuities also appear in the Hxc kernel, responsible for aforementioned orbital relaxation effects, making computation of the Fukui function within PPLB challengingHellgren and Gross (2012). Constraints on KS orbital densities arising from piecewise-linearity requirement of the physical density were recently investigatedKraisler (2025), aiming to identify and correct density-driven errors in DFT calculations. The HOMO dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems. In some cases, it was found to not even be two-point-Taylor-expandable, with preliminary analysis suggesting the divergence of Taylor coefficients stems from approximate functionals being used.
The recently introduced $N$-centered extension of ensemble density functional theorySenjean and Fromager (2018); Hodgson et al. (2021); Cernatic et al. (2022) (Nc EDFT) is an alternative exact approach to charged electronic excitations that shed a new light on the discontinuity problem. While PPLB considers an ensemble of $N$ and N1N-1 electons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra). The ensemble density and weights become independent variables of the theory, which transform the elusive energy derivative discontinuity contributions into weight-derivatives of the Hxc ensemble functional. This offers a promising way to ”recycle” local KS DFA lacking derivative discontinuities. It also provides a considerable gain in scope as it can describe any charged or neutral excitation within a unified pictureCernatic et al. (2024), and can be adapted to treat open systemsSenjean and Fromager (2020). We can benefit from progress on DFA development for EDFT in recent yearsGould and Pittalis (2019); Fromager (2020); Gould et al. (2020, 2023); Gould and Pittalis (2024); Gould et al. (2025a, b). Moreover, the introduction by the Authors of in principle exact expressions for neutrally-excited state densitiesFromager (2025) and responsesDupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
In this Letter, we derive an exact $N$-centered EDFT of Fukui functions and discuss the fundamental and practical implications of our findings in the description of second (and possibly higher)-order density-functional Hxc derivatives. We then investigate two strategies to build a practical EDFAs and study their performance on the Hubbard dimer model systemCarrascal et al. (2015) as proof of principle. First, we will show how dressing ground-state KS DFAs with a weight-dependent scaling function allows to capture the weight-derivative term. Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc kernel, with additional constraint from the ground-state uniform electron gas limit.
[…]
$n^{\bm{\xi}}({\bf r})=\left(1-\sum_{\lambda>0}\frac{N_{\lambda}}{N}\xi_{\lambda}\right)n_{\Psi_{0}}({\bf r})+\sum_{\lambda>0}\xi_{\lambda}n_{\Psi_{\lambda}}({\bf r}),$ (2)
where $n_{\Psi}(\mathbf{r}):=\langle\Psi|\hat{n}(\mathbf{r})|\Psi\rangle$, integrates to the reference (so-called central) electron number $N$ of the reference $\nu=0$ ground state. It thus generalizes the Theophilou-Gross-Oliveira-Kohn (TGOK) ensemble framework for neutral excitations Theophilou (1979); Gross et al. (1988a, b); Oliveira et al. (1988) to also describe charged excitations, and has already been explored in some detailSenjean and Fromager (2018); Deur and Fromager (2019); Hodgson et al. (2021); Cernatic et al. (2022); Cernatic and Fromager (2024); Cernatic et al. (2024), in particular considering consequences of an ensemble density $n$ and $\bm{\xi}$ being independent variables of the theory, unlike in regular (PPLB) DFT Cernatic et al. (2022). In the conventional KS formulation of Nc EDFT, the ensemble density is mapped onto a non-interacting density matrix operator $\hat{\Gamma}_{s}^{\bm{\xi}}=\left(1-\sum_{\lambda>0}\frac{N_{\lambda}}{N}\xi_{\lambda}\right)|{\Phi^{\bm{\xi}}_{0}}\rangle\langle{\Phi^{\bm{\xi}}_{0}}|+\sum_{\lambda>0}\xi_{\lambda}|{\Phi^{\bm{\xi}}_{\lambda}}\rangle\langle{\Phi^{\bm{\xi}}_{\lambda}}|$, i.e.,
[…]
$n_{\Psi_{\nu}}=\frac{N_{\nu}}{N}n^{\bm{\xi}}+\sum_{\lambda>0}(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda})\frac{\partial n^{\bm{\xi}}}{\partial\xi_{\lambda}}$.
According to the KS mapping of Eq. (3), it leads to the more explicit expression (where we dropped the $\mathbf{r}$ dependence, for the sake of compactness),
$n_{\Psi_{\nu}}=n_{\Phi^{\bm{\xi}}_{\nu}}+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\chi^{\bm{\xi}}_{s}\star\dfrac{\partial}{\partial\xi_{\lambda}}\left(v^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]\right),$ (4)
where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$ and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$. Note that the two response functions are related through the following Dyson equation,
$\big(\chi^{\bm{\xi}}\big)^{-1}=\big(\chi^{\bm{\xi}}_{s}\big)^{-1}-f^{\bm{\xi}}_{\rm Hxc},$ (5)
where $f^{\bm{\xi}}_{\rm Hxc}=f^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]$ is the Nc ensemble density-functional Hxc kernel $f^{\bm{\xi}}_{\rm Hxc}[n](\mathbf{r},\mathbf{r}^{\prime})\equiv\delta v^{\bm{\xi}}_{\rm Hxc}[n](\mathbf{r})/\delta n(\mathbf{r}^{\prime})$ evaluated at the exact ensemble density. If we further expand and rewrite Eq. (4) as follows,
$\begin{split}&\big(\chi^{\bm{\xi}}_{s}\big)^{-1}\star\left(n_{\Psi_{\nu}}-n_{\Phi^{\bm{\xi}}_{\nu}}\right)=f^{\bm{\xi}}_{\rm Hxc}\star\left(n_{\Psi_{\nu}}-\frac{N_{\nu}}{N}n^{\bm{\xi}}\right)\\ &+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}},\end{split}$ (6)
or, equivalently,
$\begin{split}\big(\chi^{\bm{\xi}}\big)^{-1}\star n_{\Psi_{\nu}}&=\big(\chi^{\bm{\xi}}_{s}\big)^{-1}n_{\Phi^{\bm{\xi}}_{\nu}}-\frac{N_{\nu}}{N}f^{\bm{\xi}}_{\rm Hxc}\star n^{\bm{\xi}}\\ &\quad+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}},\end{split}$ (7)
we finally obtain, by noticing that $\chi^{\bm{\xi}}\big(\chi^{\bm{\xi}}_{s}\big)^{-1}=1+\chi^{\bm{\xi}}\star f^{\bm{\xi}}_{\rm Hxc}$, the following exact and compact relation between the true and KS individual-state densities,
$\begin{split}n_{\Psi_{\nu}}\underset{\nu\geq 0}{=}&\big(1+\chi^{\bm{\xi}}\star f_{\rm Hxc}^{\bm{\xi}}\big)\star n_{\Phi^{\bm{\xi}}_{\nu}}-\frac{N_{\nu}}{N}\chi^{\bm{\xi}}\star f_{\rm Hxc}^{\bm{\xi}}\star n^{\bm{\xi}}\\ &+\chi^{\bm{\xi}}\star\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}.\end{split}$ (8)
Note that Eq. (8) generalizes Eq. (104) of Ref. 41 to charged excited states. As it applies universally to any $N_{\nu}$-electron ground or excited state ($N_{\nu}=N,N\pm 1,N\pm 2,\ldots$), it can be (straightforwardly) applied to the calculation of physical (interacting) Fukui functions $f_{\kappa}={\theta}_{\kappa}(n_{\Psi_{\kappa}}-n_{\Psi_{0}})$, for which $N_{\kappa}\neq N$ and ${\theta}_{\kappa}={\rm sgn}(N_{\kappa}-N)$, thus leading to the following exact relation between $f_{\kappa}$ and its non-interacting KS analog $f^{\bm{\xi}}_{s,\kappa}={\theta}_{\kappa}(n_{\Phi^{\bm{\xi}}_{\kappa}}-n_{\Phi^{\bm{\xi}}_{0}})$,
[…]
Eq. (9) is the first key result of this Letter.
Interestingly, in regular (PPLB) DFT, this relation reduces to the first term on the right-hand side (the second and third terms simply do not exist in PPLB), which is then dependent on the aforementioned discontinuous Hxc kernel for accurate predictions Hellgren and Gross (2012). It appears that, within the Nc formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq. (9). Thus, we generalize a feature that was already known for the evaluation of ionization potentials and electron affinities Cernatic et al. (2024). This equivalence can be made even more transparent by using the invariance of Eq. (9) under symmetric shifts
$f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})\rightarrow f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})+g({\bf r})+g({\bf r}^{\prime}),$ (10)
since $|\int d\mathbf{r}\,f^{\bm{\xi}}_{s,\kappa}(\mathbf{r})=\absolutevalue{N_{\kappa}-N}$ and $\int d\mathbf{r}\,\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta\left(\int d\mathbf{r}\,n^{\bm{\xi}}(\mathbf{r})\right)/\delta v(\mathbf{r}^{\prime})=\delta N/\delta v(\mathbf{r}^{\prime})$=0. This fundamental property is specific to Nc EDFT. Indeed, in PPLB, physically-relevant variations of the density lead to a change of particle number, thus breaking the gauge-invariance Hellgren and Gross (2012). Note that, within Nc EDFT, we can always make a gauge choice for a specific charged excitation $0\rightarrow\kappa]$ such that the last two terms on the right-hand side of Eq. (9) cancel out, i.e.,
$\begin{split}&\chi^{\bm{\xi}}\star\bigg[\frac{|N_{\kappa}-N|}{N}f_{\rm Hxc}^{\bm{\xi}[0\rightarrow\kappa]}\star n^{\bm{\xi}}\\ &-{\theta}_{\kappa}\sum_{\lambda>0}\Big(\delta_{\lambda\kappa}-\frac{N_{\kappa}-N}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}\bigg]=0,\end{split}$ (11)
so that Eq. (9) becomes formally identical to that of PPLB, with no approximations made. Therefore, each excitation process requires applying a specific ($\kappa$-dependent) weight-dependent symmetric shift $g^{\bm{\xi}[0\rightarrow\kappa]}({\bf r})+g^{\bm{\xi}[0\rightarrow\kappa]}({\bf r}^{\prime})=f_{\rm Hxc}^{\bm{\xi}[0\rightarrow\kappa]}(\mathbf{r},\mathbf{r}^{\prime})-f_{\rm Hxc}^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})$ to the ensemble Hxc kernel to ensure a systematic exactification of the PPLB formula, which then holds not only for ground but also for excited $N_{\kappa}$-electron states. According to Eq. (11),
$\begin{split}g^{\bm{\xi}[0\rightarrow\kappa]}&=\dfrac{1}{N_{\kappa}-N}\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\kappa}}\Big|_{n=n^{\bm{\xi}}}\\ &\quad-\dfrac{1}{N}\left(\sum_{\lambda>0}\xi_{\lambda}\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}\right)-f_{\rm Hxc}^{\bm{\xi}}\star\dfrac{n^{\bm{\xi}}}{N}.\end{split}$ (12)
Eq. (12), which is our second and central formal result, establishes the equivalence between the discontinuities in the Hxc kernel and the weight derivatives of the ensemble Hxc density-functional potential (see the first term on the right-hand side). The clarity of the Nc EDFT picture is made manifest in the fact that no assumption about the asymptotic behavior of the density needs to be invoked Hellgren and Gross (2012) to rationalize and evaluate the shift. This is relevant for analysis of periodic systems, where Fukui functions can rationalize surface reactivity Barrera et al. (2025).
Practical strategies for EDFA design – By exploring the performance of EDFAs for computing Fukui functions through Eq. (9), we will discuss what a good approximation to the ensemble Hxc functional is, gather insights on the interplay of density-dependence and weight-dependence, and particularly, how Nc EDFT could possibly recycle semi-local standard DFAs whose lack of discontinuous behavior impeded PPLB. We do so by applying the Nc formalism to the Hubbard Dimer as it allows to study all regimes of electronic correlation (controlled by the parameter $U/t$) while the Hxc functional and its derivatives can be computed exactlyCarrascal et al. (2015). Note the strategies are general, and applicable to Ab Initio systems. We will focus on charged excitations between different electronic ground states, considering an ensemble of the $N=2$-electron state and the ”neighboring” $N\pm 1$-electron states with weights $\bm{\xi}=(\xi_{-},\xi_{+})$. In all results shown, we evaluate density functionals at the exact ensemble density. The ensemble response function $\chi^{\bm{\xi}}$ will be computed from Eq. (5) using the exact KS response.
The opportunity to employ regular (ground-state) DFAs arises from noticing that Eq.(9) holds even in the zero weights limit, where the ensemble Hxc kernel, response and density become those of regular $N$-electron ground-state DFT:
$\begin{split}f_{\pm}&=\big(1+\chi\star f_{\rm Hxc}\big)\star f_{s,\pm}-\;\chi\star f_{\rm Hxc}\star\frac{n_{0}}{N}\\ &\quad\pm\;\chi\star\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n_{0}}^{\bm{\xi}=0},\end{split}$ (13)
$f_{s,+}(\mathbf{r})$ is the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state and $f_{s,-}(\mathbf{r})$ the highest occupied one (HOMO). Note $n_{0}=n^{\bm{\xi}=0}=n_{\Psi_{0}}$.
This is an astounding simplification: the only piece needed to compute Fukui functions aside from a ground-state functional without any approximation is the weight-derivatives the ensemble Hxc potential evaluated for null weights. In practice, one does not need to build an ensemble DFA covering the whole range of valid weights to obtain accurate results, but merely to model the weight dependence of the potential in the immediate vicinity of the familiar $N$-electron ground-state limit.
To that aim, it is appealing to ”dress” a regular KS-DFA with a weight-dependent scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n]$ (14)
where $s(\bm{\xi}=0)=1$ by construction. The weight-derivatives would then be approximated as
$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\lambda}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}]$ (15)
This takes full advantage of the 0-weight limit mindset, $s(\bm{\xi})$ only accounting for the ensemble behavior in the vicinity of $\bm{\xi}=0$, and benefits from all the known exact constraints built in the ground-state functional provided the scaling function is chosen in a way that complies with them. In particular, this choice complies with exact uniform coordinate scaling constraint valid both for regular KS DFA and for ensembles Nagy (1995); Scott et al. (2024). Known limits of the ensemble Hxc functional can be used to build $s(\bm{\xi})$, as we will explore in the following. Throughout this work, we will use the exact 2-electrons Hubbard ground-state Hxc functional in pair with scaling functions.
Low correlation regime – Orbital-based approximations were used extensively in EDFT to bypass development of explicit (weight-dependent) density functionals, the simplest being an Hartree-Exchange approximation to the ensemble energyYang et al. (2014); Pribram-Jones et al. (2014); Yang et al. (2017); Gould et al. (2018); Sagredo and Burke (2018); Deur et al. (2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020). When the correlation is not too strong, one can expect such an approach to do well, which we put to the test by considering the exact exchange ensemble functional (EEXX) of the Hubbard Dimer. We recall it is obtained as the first order term in the Taylor expansion in $U$ of the universal functional,
$E^{\bm{\xi}}_{\rm Hx}[n]=U\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}$ (16)
and yields all quantities in Eq. (9) by differentiation of $E^{\bm{\xi}}_{\rm Hx}[n]$. All results are computed in the 0-weight limit.
Figure 1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset). Full and dashed lines are $f_{+}$ and $f_{-}$ calculations respectively. Exact results (yellow lines) are compared to EEXX (blue lines) and to the neglect of weight-derivative contributions (grey lines).
In the weak correlation regime $U/t=0.75$, the EEXX functional closely reproduces exact results as seen on Fig. 1, with the effect of neglecting Hxc weight derivatives shown in grey. The improvement is clear, making this first application of EEXX to compute densities promising.
These encouraging results suggest using the scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hxc}[n]\,,\qquad s_{\rm Hx}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}{v_{\rm Hx}[n^{\bm{\xi}}]}$ (17)
Note the ensemble density is fed to the ground-state potential. It is important to define the scaling in this way to ensure evaluating the weight-derivative of the scaling function (in the 0-weight limit) does not require knowing the weight-derivative of the ensemble density. While using this scaling function misses correlation at the ensemble level, note some measure of it is still brought into the approximate potential through the ground-state Hxc potential (see Eq. (17)). For $U/t=0.75$, there is negligible difference between the EEXX scaling method (not shown) and full EEXX DFA, which could be expected as correlation barely contributes in this regime.
Figure 2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to EEXX scaling (red) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Intermediate correlation regime – We expect this to change by turning to higher (but moderate) correlation $U/t=1.5$, results being shown on Fig. 2. For external potentials below $\Delta v=2U/t$, where the correlation energy becomes important, we see EEXX deviates from the exact result. It still improves substantially over the neglect of weight-derivative terms for $f_{-}$, but not for $f_{+}$: the weight-derivative terms only bring a small correction here, which is overestimated in EEXX. Turning to the scaling strategy, we ask: Is ground-state correlation a reasonable approximation of its ensemble analog? Results show dressing the ground-state correlation potential with a purely Hx ensemble scaling brings a modest improvement over full EEXX for $f_{-}$, but prediction of $f_{+}$ actually worsens.
These results show it is necessary to go beyond EEXX when correlation takes over. Moreover, we have to understand if the failure of Hx scaling function comes from an inherent discrepancy between ground-state and ensemble correlation, making the weight-dependent dressing strategy impractical, or if it can be fixed. To deal with situations where EEXX fails, second order (Görling-LeviGörling and Levy (1994)) perturbation theory has been applied to ensemble DFT, yielding an orbital-dependent DFA beyond EEXXYang (2021). We transpose this in the Hubbard dimer through second-order Taylor expansion in $U$:
$E^{\bm{\xi}}_{\rm c}[n]\overset{\rm PT2}{\approx}\frac{U^{2}}{2}\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}$ (18)
Compared to EEXX-only on Fig. 3, it shows much better agreement with exact Fukui functions.
Figure 3: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to PT2 (green) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Moreover, comparison of functional forms of EEXX and PT2 shows where our Hx-only scaling is failing: The ratio of PT2 ensemble and ground-state correlation potentials is an explicit function of the weights and of the ensemble density:
$s_{\rm c,PT2}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm c,PT2}[n^{\bm{\xi}}]}{v_{\rm c,PT2}[n^{\bm{\xi}}]}=s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})$ (19)
while in the Hx ratio Eq. (17) all dependence on the ensemble density simplifies out, leaving only explicit weight dependences. This disparity in complexity for the Hx and correlation part suggests using a different scaling function for each, which reads at the PT2 level:
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})\;v_{\rm c}[n]$ (20)
with exact ensemble density dependence shown explicitly. Note if the correlation term was neglected all-together, leaving only the first term, the potential would be equal to that of EEXX (in the 0-weight limit). But at the PT2 level, it dresses the exact ground-state correlation with an ensemble-density dependence. The question now becomes: Does this double scaling at the PT2 level improve over the full PT2 EDFA? On Fig. 4, we compare their performance in computing $f_{+}$ for $U/t$ up to 2.5. We can see that the PT2 scaling accurately matches exact results even above $U=1.5$, when both EEXX and PT2 fail. This shows the scaling function strategy holds promise, but it is crucial to treat the Hx and correlation parts separately.
Figure 4: Fukui function $f_{+}$ in the low correlation regime for null ensemble weights (parameters in the red inset). For increasing $U$, we compare our EEXX EDFA (blue) to PT2 EDFA (green) and the PT2 scaling strategy (red). Exact results (orange) and neglect of weight-derivatives (grey) are plotted for reference.
Strong correlation – Learning from high- and low-correlation limits of the EDFT proved to be an outstanding approach for excitation energies in difficult situationsGould et al. (2023), with a versatile EDFA obtained from interpolating between the two regimes. We put it to the test for Fukui functions by building a Padé interpolation between the PT2 (low $U$) and $U/t\rightarrow\infty$ limits of the ensemble Hxc functional,
$E_{\rm Hxc}^{\bm{\xi}}[n]=a^{\bm{\xi}}[n]\,U+\frac{b^{\bm{\xi}}[n]\,U^{2}}{1+c^{\bm{\xi}}[n]\,U}$ (21)
where
$\begin{split}a^{\bm{\xi}}[n]&=\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}\\ b^{\bm{\xi}}[n]&=\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}\\ c^{\bm{\xi}}[n]&=\frac{b^{\bm{\xi}}[n]}{\gamma^{\bm{\xi}}[n]-a^{\bm{\xi}}[n]}\end{split}$ (22)
and the strong-correlation-dominant term
$\gamma^{\bm{\xi}}[n]=\lim_{U/t\rightarrow\infty}E_{\rm Hxc}^{\bm{\xi}}[n]={\rm sup}\Big\{\xi_{+},\frac{\xi_{+}-\xi_{-}}{2}+\big|n-1\big|\Big\}$ (23)
Figure 5: Fukui functions in strong correlation regime for non-zero weights (parameters in the red inset). We compare the Padé EDFA (green) to EEXX prediction (blue). Exact results (orange) and neglect of the weight-derivative terms (grey) are plotted for reference.
This EDFA performs poorly (not shown) as its discontinuous behavior (through the sup function) and density-derivative discontinuity in |n1||n-1| means it lacks cross-derivatives (in $n$ and $\xi_{\pm}$) contributions to $\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n^{\bm{\xi}}}$ that are important even in the high correlation regime. Even if Eq. (23) is formally exact in the strictly-correlated regime ($U\rightarrow\infty$), it is not representative of the strongly-correlated regime (typically $U\in[5,10]$). To remedy this, the sharp sup function and absolute value were each replaced by a smoothing function whose stiffness parameters increase with $U$, thus yielding cross derivative terms while still being compatible with the exact $U/t\rightarrow\infty$ limit. To set the stiffness parameters in a way consistent with what could be done for a real system, we only considered the symmetric case of 2-electrons Hubbard dimer, i.e. $\Delta v=0$ and $\bm{\xi}=0$, which is the analog of the uniform electron gas in this context. The stiffness parameters were set so that our Padé approximation reproduces best the (ground-state) Hxc kernel. The obtained EDFA is in principle applicable for any choice of weights, even if it was smoothed by learning from the ground-state symmetric case. It is thus most interesting to test its performance for non-zero weights.
[…]
This trend carries over to higher correlation, as shown on Fig. 6 for $U/t=10$ and $\xi=0.2$, which is remarkable as the stifness parameters smoothing the $U\rightarrow\infty$ discontinuities in $\gamma^{\bm{\xi}}[n]$ where fixed using only the ground-state, symmetric limit of the Hxc functional.
Figure 6: Fukui functions in very strong correlation regime for non-zero ensemble weights (parameters in the red inset). See Fig. 5 for labels.
Summary and outlook– After introducing the (in principle) exact expression of Fukui functions within Nc EDFT, we discussed the equivalence between weight-derivatives of the Hxc potential and derivative discontinuities in the Hxc kernel. The exact relations introduced by the Authors between individual true physical and non-interacting KS densitiesFromager (2025) and responsesDupuy and Fromager (2025) are thus a promising framework to access properties beyond excitation energies within EDFT.
By considering the conventional $N$-electron ground-state DFT limit of the theory, we showed how dressing the regular (ground-state) Hxc functional with a weight-dependent scaling function allows to recycle practical (semi-)local DFAs while accurately accounting for the ensemble Hxc potential weight-derivatives term, as illustrated on Fukui functions of the Hubbard Dimer. It was instrumental to the performance of this strategy that separate scaling functions would be applied to the Hx and correlation parts of the ground-state potential, as their ensemble analogs had starkly different dependence on ensemble weights and density. This is in line with analysis of low- and high-correlation limits of EDFTGould et al. (2023), predicting a dominant dependence on the ensemble density rather than on the weights in the latter case. When used properly, the scaling function approach extended the domain of applicability of its parent approximation by exploiting an accurate ground-state functional.
We also showed building a full ensemble DFA by interpolating between known low- and high-correlation limits performed very well, provided one paid attention to avoid discontinuous behavior in the functional. Smoothing over these discontinuities by learning from the symmetric ground state proved very effective, echoing recent work on the uniform electron gas as a basis to develop excited state density functionalsLoos (2025). These findings will foster DFA development in realistic systems. As additional strategies, one could also include exact constraints satisfied by ensembles in the design of scaling functions, such as the invariance of individual-state physical densities with respect to weights (see Eq. (8)), which is currently explored. As our formalism applies to any (neutral or charged) excitation, it can bring insights into processes that PPLB could not grasp, notably excited state reactivity through the relevant Fukui functionsAn et al. (2024). This is left for future work.
Acknowledgements - This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).

### v2

@@
 Abstract
-We introduce an in-principle exact working equation to compute Fukui functions within $N$-centered ensemble DFT.
-It avoids the kernel derivative discontinuity problem of DFT for fractional number of electrons, whose contribution is recovered through weight-derivatives of the ensemble density functional potential.
-We explore practical strategies to compute its contribution by recycling ground-state density-functional approximations dressed with a weight-dependent scaling function.
-We also show interpolating between known limits of the ensemble functional and learning from uniform density profiles is a very effective strategy.
+An in-principle exact working equation to compute electronic affinity and ionization Fukui functions is derived within the $N$-centered (Nc) ensemble extension of density functional theory (DFT).
+It circumvents the kernel derivative discontinuity problem of DFT for fractional electron numbers, whose contribution is recovered through weight derivatives of the ensemble density functional potential.
+Thus, it allows for the design of alternative and effective approximations, such as the weight-dependent scaling of regular functionals or the interpolation between known limits of Nc ensembles.
 Introduction – No matter how much we succeed in making accurate computer simulations of chemical reactions routine, there will always be a need for chemical descriptors of reactivity.
@@
 Further analysis of increasingly local responses of the electronic energy to changes in the external potential and number of electrons bloomed into the field of Conceptual Density Functional Theory Geerlings et al.
-(2003); Geerlings and De Proft (2008) (CDFT).
+(2003); Geerlings and Proft (2008) (CDFT).
 In turn, their computation from corresponding functional derivatives of the (to-be-approximated) universal functional proved instrumental in understanding what information practical Density Functional Approximations (DFA) should incorporate: By focusing efforts on reproducing more relevant properties, DFA design can (hopefully) find shortcuts on the meandering path towards the exact functional Gould (2017).
@@
 It offers a (sometimes critical Bartolotti and Ayers (2005)) improvement over frontier molecular orbital theory Fukui et al.
-(1952, 1954); Fukui (1982) by inclusion of orbital relaxation effects.
+(1952, 2006); Fukui (1982) by inclusion of orbital relaxation effects.
 Interpretation and computation of derivatives with respect to the number of electrons analytically, i.e.
-connecting them to the $N$-electron Kohn Sham system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and NN-1 electron systemsYang et al.
+connecting them to the $N$-electron Kohn Sham (KS) system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and N1N-1 electron systems Yang et al.
 (2012).
@@
 Constraints on KS orbital densities arising from piecewise-linearity requirement of the physical density were recently investigated Kraisler (2025), aiming to identify and correct density-driven errors in DFT calculations.
-The HOMO dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems.
+The highest occupied molecular orbital (HOMO) dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems.
 In some cases, it was found to not even be two-point-Taylor-expandable, with preliminary analysis suggesting the divergence of Taylor coefficients stems from approximate functionals being used.
@@
 (2022) (Nc EDFT) is an alternative exact approach to charged electronic excitations that shed a new light on the discontinuity problem.
-While PPLB considers an ensemble of $N$ and N1N-1 electons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra).
+While PPLB considers an ensemble of $N$ and N1N-1 electrons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra).
 The ensemble density and weights become independent variables of the theory, which transform the elusive energy derivative discontinuity contributions into weight-derivatives of the Hxc ensemble functional.
-This offers a promising way to ”recycle” local KS DFA lacking derivative discontinuities.
+This offers a promising way to ”recycle” (semi-) local DFAs lacking derivative discontinuities Loos and Fromager (2020).
 It also provides a considerable gain in scope as it can describe any charged or neutral excitation within a unified picture Cernatic et al.
 (2024), and can be adapted to treat open systems Senjean and Fromager (2020).
-We can benefit from progress on DFA development for EDFT in recent yearsGould and Pittalis (2019); Fromager (2020); Gould et al.
+The theory can also benefit from progress made in recent years in ensemble DFA development Gould and Pittalis (2019); Fromager (2020); Gould et al.
 (2020, 2023); Gould and Pittalis (2024); Gould et al.
-(2025a, b).
-Moreover, the introduction by the Authors of in principle exact expressions for neutrally-excited state densitiesFromager (2025) and responsesDupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
+(2026, 2025).
+Moreover, the very recent derivation by the authors of exact expressions for neutrally-excited state densities Fromager (2025) and their responses Dupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
 In this Letter, we derive an exact $N$-centered EDFT of Fukui functions and discuss the fundamental and practical implications of our findings in the description of second (and possibly higher)-order density-functional Hxc derivatives.
-We then investigate two strategies to build a practical EDFAs and study their performance on the Hubbard dimer model systemCarrascal et al.
+We then investigate two strategies to build a practical EDFA and study their performance on the Hubbard dimer model system Carrascal et al.
 (2015) as proof of principle.
-First, we will show how dressing ground-state KS DFAs with a weight-dependent scaling function allows to capture the weight-derivative term.
-Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc kernel, with additional constraint from the ground-state uniform electron gas limit.
+First, we will show how dressing the regular ground-state Hxc functional with a weight-dependent scaling function allows to capture the weight-derivative term.
+Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc ensemble density functional.
 […]
@@
 […]
-$n_{\Psi_{\nu}}=\frac{N_{\nu}}{N}n^{\bm{\xi}}+\sum_{\lambda>0}(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda})\frac{\partial n^{\bm{\xi}}}{\partial\xi_{\lambda}}$.
-According to the KS mapping of Eq. (3), it leads to the more explicit expression (where we dropped the $\mathbf{r}$ dependence, for the sake of compactness),
 $n_{\Psi_{\nu}}=n_{\Phi^{\bm{\xi}}_{\nu}}+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\chi^{\bm{\xi}}_{s}\star\dfrac{\partial}{\partial\xi_{\lambda}}\left(v^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]\right),$ (4)
-where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$ and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$.
+where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$, and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$.
 Note that the two response functions are related through the following Dyson equation,
@@
 Interestingly, in regular (PPLB) DFT, this relation reduces to the first term on the right-hand side (the second and third terms simply do not exist in PPLB), which is then dependent on the aforementioned discontinuous Hxc kernel for accurate predictions Hellgren and Gross (2012).
-It appears that, within the Nc formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq.
+It appears that, within the Nc EDFT formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq.
 (9).
@@
 This equivalence can be made even more transparent by using the invariance of Eq.
-(9) under symmetric shifts
+(9) under symmetric shifts in the ensemble Hxc kernel,
 $f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})\rightarrow f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})+g({\bf r})+g({\bf r}^{\prime}),$ (10)
@@
 This is relevant for analysis of periodic systems, where Fukui functions can rationalize surface reactivity Barrera et al.
+(2025), for example.
+Practical strategies for Nc EDFA design – The opportunity to employ regular (ground-state) DFAs in this context arises from noticing that Eq.
+(9) holds even in the zero-weight limit of the theory, where the ensemble Hxc kernel, response function and density become those of regular $N$-electron ground-state DFT.
+If we focus more specifically on the ground-state theory of a single electron addition ($\kappa\equiv+$) or removal ($\kappa\equiv$-), it reads
+$\begin{split}f_{\pm}&=\big(1+\chi\star f_{\rm Hxc}\big)\star f_{s,\pm}-\;\chi\star f_{\rm Hxc}\star\frac{n_{0}}{N}\\ &\quad\pm\;\chi\star\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n_{0}}^{\bm{\xi}=0},\end{split}$ (13)
+$f_{s,+}$ being the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state, $f_{s,-}$ the HOMO, and $n_{0}:=n^{\bm{\xi}=0}=n_{\Psi_{0}}$.
+This is an astounding simplification where the only piece needed (aside from the regular $N$-electron Hxc kernel) to compute Fukui functions, in principle exactly, is the weight-derivatives of the Nc ensemble density-functional Hxc potential evaluated at $\bm{\xi}\equiv\{\xi_{+},\xi_{-}\}=0$.
+In practice, one does not need to build an EDFA covering the whole range of positive weight values to obtain accurate results, but merely to model the weight dependence of the Hxc potential in the immediate vicinity of the familiar $N$-electron ground-state limit.
+To that aim, it is appealing to “dress” any regular DFA with a weight-dependent scaling function,
+$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n],$ (14)
+where $s(\bm{\xi}=0)=1$, by construction.
+In principle, one could use Hxc energies ratio to define $s(\bm{\xi})$, but as illustrated in the supplemental material, their weight-dependence differ widely from that of the potentials, making them unsuited for the task.
+Unlike a bare (weight-independent) DFA, the above EDFA will enable the description of weight-derivatives, as follows,
+$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\pm}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}].$ (15)
+While being useful, it finds its theoretical justification in the uniform coordinate scaling exact constraints that both DFA and EDFA are expected to fulfill Nagy (1995); Scott et al.
 (2024).
-Practical strategies for EDFA design – By exploring the performance of EDFAs for computing Fukui functions through Eq.
-(9), we will discuss what a good approximation to the ensemble Hxc functional is, gather insights on the interplay of density-dependence and weight-dependence, and particularly, how Nc EDFT could possibly recycle semi-local standard DFAs whose lack of discontinuous behavior impeded PPLB.
-We do so by applying the Nc formalism to the Hubbard Dimer as it allows to study all regimes of electronic correlation (controlled by the parameter $U/t$) while the Hxc functional and its derivatives can be computed exactlyCarrascal et al.
+Known limits of the exact ensemble Hxc functional can be used to build $s(\bm{\xi})$, for example in the weakly and strictly correlated regimes Gould et al.
 (2023).
-Note the strategies are general, and applicable to Ab Initio systems.
-We will focus on charged excitations between different electronic ground states, considering an ensemble of the $N=2$-electron state and the ”neighboring” $N\pm 1$-electron states with weights $\bm{\xi}=(\xi_{-},\xi_{+})$.
-In all results shown, we evaluate density functionals at the exact ensemble density.
-The ensemble response function $\chi^{\bm{\xi}}$ will be computed from Eq.
-(5) using the exact KS response.
-The opportunity to employ regular (ground-state) DFAs arises from noticing that Eq.(9) holds even in the zero weights limit, where the ensemble Hxc kernel, response and density become those of regular $N$-electron ground-state DFT:
-$\begin{split}f_{\pm}&=\big(1+\chi\star f_{\rm Hxc}\big)\star f_{s,\pm}-\;\chi\star f_{\rm Hxc}\star\frac{n_{0}}{N}\\ &\quad\pm\;\chi\star\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n_{0}}^{\bm{\xi}=0},\end{split}$ (13)
-$f_{s,+}(\mathbf{r})$ is the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state and $f_{s,-}(\mathbf{r})$ the highest occupied one (HOMO).
-Note $n_{0}=n^{\bm{\xi}=0}=n_{\Psi_{0}}$.
-This is an astounding simplification: the only piece needed to compute Fukui functions aside from a ground-state functional without any approximation is the weight-derivatives the ensemble Hxc potential evaluated for null weights.
-In practice, one does not need to build an ensemble DFA covering the whole range of valid weights to obtain accurate results, but merely to model the weight dependence of the potential in the immediate vicinity of the familiar $N$-electron ground-state limit.
-To that aim, it is appealing to ”dress” a regular KS-DFA with a weight-dependent scaling function
-$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n]$ (14)
-where $s(\bm{\xi}=0)=1$ by construction. The weight-derivatives would then be approximated as
-$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\lambda}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}]$ (15)
-This takes full advantage of the 0-weight limit mindset, $s(\bm{\xi})$ only accounting for the ensemble behavior in the vicinity of $\bm{\xi}=0$, and benefits from all the known exact constraints built in the ground-state functional provided the scaling function is chosen in a way that complies with them.
-In particular, this choice complies with exact uniform coordinate scaling constraint valid both for regular KS DFA and for ensembles Nagy (1995); Scott et al.
-(2024).
-Known limits of the ensemble Hxc functional can be used to build $s(\bm{\xi})$, as we will explore in the following.
-Throughout this work, we will use the exact 2-electrons Hubbard ground-state Hxc functional in pair with scaling functions.
-Low correlation regime – Orbital-based approximations were used extensively in EDFT to bypass development of explicit (weight-dependent) density functionals, the simplest being an Hartree-Exchange approximation to the ensemble energyYang et al.
+Considering the former, a sensible pick would consist in combining approximations based on the ensemble exact exchange (EEXX), which has been used extensively to compute excitation energies Yang et al.
 (2014); Pribram-Jones et al.
@@
 (2018); Sagredo and Burke (2018); Deur et al.
-(2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020).
-When the correlation is not too strong, one can expect such an approach to do well, which we put to the test by considering the exact exchange ensemble functional (EEXX) of the Hubbard Dimer.
-We recall it is obtained as the first order term in the Taylor expansion in $U$ of the universal functional,
-$E^{\bm{\xi}}_{\rm Hx}[n]=U\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}$ (16)
-and yields all quantities in Eq. (9) by differentiation of $E^{\bm{\xi}}_{\rm Hx}[n]$. All results are computed in the 0-weight limit.
-Figure 1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset).
+(2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020), with a perturbative treatment of electron correlation at the ensemble level, as proposed in Ref.
+59.
+A priori, this strategy cannot tackle strongly correlated systems.
+While the scaling approach of Eq.
+(14) may improve on this thanks to the partial account of correlation through the ground-state functional, a more global strategy that interpolates between extreme limits is needed to ensure that EDFA can describe all correlation regimes within the ensemble.
+This proved to be an outstanding approach for excitation energies in difficult situations Gould et al.
+(2023), for example.
+We show in the following how these different strategies can be implemented successfully for the computation of Fukui functions.
+Illustrative example – As a proof of principle, we apply the present Nc EDFT formalism to the simple but nontrivial asymmetric Hubbard dimer Carrascal et al.
+(2015, 2016); Fromager (2020).
+The model allows to study all regimes of electronic correlation, by tuning the on-site (atomic) electronic repulsion to the electronic hopping (which is analogous to the kinetic energy strength) ratio $U/t$, for a given difference $\Delta v$ in external potential between the two sites, while the ensemble Hxc functional and its derivatives can be computed exactly.
+We focus on charged excitations between different electronic ground states, as described in Eq.
+(13), with the central number of electrons set to $N=2$.
+Exact ensemble densities are systematically used so that we can focus on functional-driven (and, in particular, weight-driven) errors.
+We start with the following EEXX-scaled functional, $v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})v_{\rm Hxc}[n]$, where the scaling function $s_{\rm Hx}(\bm{\xi})={v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}/{v_{\rm Hx}[n^{\bm{\xi}}]}$ relies on the (analytically solvable in the model) EEXX scheme.
+Note that, in the targeted zero-weight limit, the weight derivative of the EEXX scaling function does not involve the response of the ensemble density to weight variations, by construction.
+While reducing to regular EEXX when electron correlation is negligible, the above EDFA can in principle capture some correlations through the ground-state correlation potential contribution $s_{\rm Hx}(\bm{\xi})v_{\rm c}[n]$.
+Figure 1: Performance of the EEXX-scaled EDFA in evaluating Fukui functions (as functions of the external asymmetry potential $\Delta v$) from the zero-weight limit of Nc EDFT in the moderately correlated $U/t=1.5$ regime of the Hubbard dimer.
+Results are shown in Fig.
+1 in the moderately correlated $U/t=1.5$ regime.
+While EEXX deviates from the exact result, as expected, it still improves substantially over the complete neglect of weight derivatives in the evaluation of the ionization Fukui function $f_{-}$, but not for the affinity one ($f_{+}$).
+In the latter case, ensemble correlation effects mostly compensate the EEXX ones.
+Unfortunately, the EEXX-scaled functional worsens the prediction of $f_{+}$, while bringing a modest improvement over EEXX for $f_{-}$.
+Therefore, it seems essential to correlate the scaling function.
+For that purpose, the extension to ensembles Yang (2021) of Görling–Levy perturbation theory Görling and Levy (1993, 1994) through second order (PT2) can be used.
+As shown in the supplementary material, unlike in EEXX, the density-functional correlation scaling function $v^{\bm{\xi}}_{\rm c,PT2}[n]/v_{\rm c,PT2}[n]=:s_{\rm c,PT2}[\bm{\xi},n]$ is both weight- and density-dependent in the model, thus suggesting the following refinement of the ansatz introduced in Eq.
+(14),
+$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}[\bm{\xi},n^{\bm{\xi}}]\;v_{\rm c}[n],$ (16)
+with a separate scaling for Hx and correlation potentials.
+While PT2 (applied on top of EEXX) fixes the flaw of EEXX mentioned previously, when $U/t=1.5$, the PT2-scaled ansatz of Eq.
+(16) has the advantage of extending the applicability of PT2 (through the scaling function) to stronger correlation regimes, with a remarkably high accuracy, as illustrated in Fig.
+2 for the affinity Fukui function.
+A deeper analysis of the correlation potential, which illustrates how the scaling approach corrects erroneous behaviors of PT2 for strongly asymmetric dimers, is provided in the supplemental material.
+It clearly shows that the scaling function strategy holds promises, but it is crucial to treat Hx and correlation parts separately.
+Figure 2: Performance of the PT2-scaled EDFA in evaluating the affinity Fukui function (as a function of the interaction strength $U$) up to the stronger correlation $U/t=2.5$ regime of the asymmetric ($\Delta v/t=3.0$) Hubbard dimer.
+For completeness, we also extended to charged excitations the ”interpolation between extreme limits” strategy of Ref.
+37.
+In the Hubbard dimer, it translates into the following Padé approximant (see the supplementary material for the complete expression) to the ensemble density-functional Hxc energy,
+$E_{\rm Hxc}^{\bm{\xi}}[n]\approx a^{\bm{\xi}}[n]{U}+\frac{b^{\bm{\xi}}[n]U^{2}}{1+\frac{b^{\bm{\xi}}[n]}{\gamma^{\bm{\xi}}[n]-a^{\bm{\xi}}[n]}U},$ (17)
+which is exact in both weakly ($U/t\rightarrow 0$) and strictly correlated ($U/t\rightarrow+\infty$) limits, by construction, and from which the Fukui functions can be evaluated through density and weight derivatives.
+To be applicable (and numerically stable) to intermediate correlation regimes, a smoothed version of the coefficient $\gamma^{\bm{\xi}}[n]$ in Eq.
+(17) has been designed such that the Hxc kernel is reproduced at all interaction strengths in the pure ground-state ($\bm{\xi}=0$) and symmetric ($\Delta v=0$) case.
+The latter is the analog of the uniform electron gas in this context, echoing recent work to develop (semi-) local density functionals for excited states Loos (2025).
+Figure 3: Same as Fig. 1 for the Padé approximant-based EDFA in the strongly correlated $U/t=5$ regime and away from the zero-weight limit ($\xi_{+}=\xi_{-}=0.2$).
+The robustness of the smoothed Padé approximant away from the zero-weight limit it learned from (we take $\xi_{+}=\xi_{-}=\xi=0.2$) is illustrated in Fig.
+3, where it successfully applies to the strongly correlated $U/t=5$ dimer in all asymmetry regimes, unlike EEXX, as expected.
+In summary, an alternative ($N$-centered) EDFT of Fukui functions has been derived, thus revealing that important but challenging higher-order Hxc derivative discontinuities can be tackled, in principle exactly, through ensemble weight derivative corrections to regular DFT.
+As a result, significant progress made recently in the understanding and application of EDFT Gould et al.
+(2026, 2025) can now be exploited for computing accurately charged excited-state properties beyond the energy, as illustrated in this Letter.
+This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).
+Supplemental Material
+Appendix S-I Analytical expressions for the Hubbard dimer
+S-I.1 Exact expressions
+Model – We consider a model diatomic system built from the Hubbard dimer, with Hamiltonian
+$\hat{H}_{e}=-t\sum_{\sigma}\big(\hat{a}_{0\sigma}^{\dagger}\hat{a}_{1\sigma}+\hat{a}_{1\sigma}^{\dagger}\hat{a}_{0\sigma}\big)+U\sum_{i=0}^{1}\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}+\frac{\Delta v}{2}\big(\hat{n}_{1}-\hat{n}_{0}\big)$ (S1)
+[…]
+$H_{N}=\begin{bmatrix}U-\Delta v&-\sqrt{2}\,t&0\\ -\sqrt{2}\,t&0&-\sqrt{2}\,t\\ 0&-\sqrt{2}\,t&U+\Delta v\end{bmatrix}$ (S2)
+the exact energies being analytically given by the third-order polynomial equationCarrascal et al. (2015, 2016); Deur et al. (2017):
+$E_{N,i}^{3}=-4t^{2}U+(4t^{2}-U^{2}+\Delta v^{2})E_{N,i}+2U\,E_{N,i}^{2}$ (S3)
+In particular, the $N=2$-electron ground state energy, that is the lowest solution to Eq. (S3), readsCarrascal et al. (2015, 2016)
+$E_{N}(\Delta v)=\frac{4t}{3}\bigg(u-w\sin\left(\theta+\frac{\pi}{6}\right)\bigg)$ (S4)
+The angle $\theta$ is defined as
+$\cos(3\theta)=\frac{u}{w^{3}}\Big(9\big(\nu^{2}-1/2\big)-u^{2}\Big)$ (S5)
+with $u=U/2t$, $w=\sqrt{3(1+\nu^{2})+u^{2}}$ and $\nu=\Delta v/2t$.
+Ensemble density and Fukui functions from the external potential – The ensemble is comprised of the 1-to-3 electron(s) ground-states, which all separately satisfy the variational principle
+$E_{N\pm p}[v]=\underset{n_{N}\pm p}{\rm min}\Big\{F[n_{N\pm p}]+\int d{\bf r}\;v({\bf r})\;n_{N\pm p}({\bf r})\Big\}$ (S6)
+with $F[n_{N\pm p}]$ the universal Hohenberg-Kohn functional. By the means of a Legendre-Fenchel transform, we can obtain an expression for the latterDeur et al. (2018):
+$F[n_{N\pm p}]=\underset{v}{\rm sup}\Big\{E_{N\pm p}[v]-\int d{\bf r}\;v({\bf r})\;n_{N\pm p}({\bf r})\Big\}$ (S7)
+from which we can deduce the link between density and energy-derivative
+$n_{N\pm p}({\bf r})=\frac{\delta E_{N\pm p}[v]}{\delta v({\bf r})}$ (S8)
+In Hubbard dimer, the Site Occupation Functional Theory (SOFT) equivalent is
+$F_{N\pm p}(\Delta n_{N\pm p})=\underset{\Delta v}{\rm sup}\Big\{E_{N\pm p}[\Delta v]-\frac{\Delta v}{2}\;\Delta n_{N\pm p}\Big\}$ (S9)
+with $\Delta n_{N\pm p}$ the density difference between site 1 and site 0.
+As the sum of density on both sites always equals $N\pm p$, we can use the density on site 0, noted $n_{N\pm p}$, as the basic variable
+$\frac{\Delta n_{N\pm p}}{2}=\frac{N\pm p}{2}-n_{N\pm p}=\frac{\partial E_{N\pm p}[\Delta v]}{\partial\Delta v}$ (S10)
+To obtain the ground-state energy of the 1-electron system we note it lacks electronic correlation entirely, and is thus given by the lowest orbital energy in the non-interacting 2-electron systemSenjean and Fromager (2018) with external potential $\Delta v$
+$E_{N_{-}}(\Delta v)=-\sqrt{t^{2}+(\Delta v/2)^{2}}$ (S11)
+while hole-particle symmetrySenjean and Fromager (2018) allows to obtain the 3-electrons ground-state energy as
+$E_{N_{+}}(\Delta v)=U-\sqrt{t^{2}+(\Delta v/2)^{2}}$ (S12)
+Inserting Eqs. (S4,S11,S12) in Eq. (S10) yields
+$\left\{\begin{split}\;&n_{N_{-}}(\Delta v)=\frac{1}{2}+\frac{1}{4}\frac{\Delta v}{\sqrt{t^{2}+(\Delta v/2})^{2}}\\ &n_{N}(\nu)=1+\frac{2\nu}{w}\sin\left(\theta+\frac{\pi}{6}\right)+\Theta\cos\left(\theta+\frac{\pi}{6}\right)\\ &n_{N_{+}}(\Delta v)=\frac{3}{2}+\frac{1}{4}\frac{\Delta v}{\sqrt{t^{2}+(\Delta v/2})^{2}}\end{split}\right$.
+(S13)
+where
+$\Theta=-\frac{4\nu}{w\sin(3\theta)}\left(\frac{u}{w}-\frac{1}{2}\cos(3\theta)\right)$ (S14)
+From Eqs. (S13), computation of the ensemble density is straightforward
+$n^{\bm{\xi}}(\Delta v)=\xi_{+}\,n_{N_{+}}(\Delta v)+\xi_{-}\,n_{N_{-}}(\Delta v)+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)n_{N}(\Delta v)$ (S15)
+as is the computation of Fukui functions. The latter are defined as the difference between densities on site 0 for neighboring $N$ and $N\pm 1$-electron ground states
+$\left\{\begin{split}\;&f_{-}(\Delta v)=n_{N}(\Delta v)-n_{N_{-}}(\Delta v)\\ &f_{+}(\Delta v)=n_{N_{+}}(\Delta v)-n_{N}(\Delta v)\end{split}\right$. (S16)
+KS Fukui functions and KS ensemble response from the KS potential – Each KS ground state within the ensemble satisfies the variational principle
+$\mathcal{E}_{N\pm p}(\Delta v^{\bm{\xi}}_{s})={\rm min}_{\Delta n_{N\pm p}}\left\{T_{s}(\Delta n)+\Delta n_{N\pm p}\,\frac{\Delta v^{\bm{\xi}}_{s}}{2}\right\}$ (S17)
+The minimizing, weight-dependent KS density differences are written $\Delta n^{\bm{\xi}}_{s,N\pm p}$. Differentiating the above expression with respect to the ensemble KS potential yields
+$\frac{\partial\mathcal{E}_{N\pm p}}{\partial\Delta v^{\bm{\xi}}_{s}}(\Delta n^{\bm{\xi}}_{s,N\pm p})=\frac{\Delta n^{\bm{\xi}}_{s,N\pm p}}{2}$ (S18)
+To obtain the total KS energies, we note KS orbital energies in the Hubbard dimer are
+$\begin{split}\epsilon_{0,1}(\Delta v^{\bm{\xi}}_{s})&=\pm\,\epsilon_{\rm H}(\Delta v^{\bm{\xi}}_{s})\\ \epsilon_{\rm H}(\Delta v^{\bm{\xi}}_{s})&=-\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2)^{2}}\end{split}$ (S19)
+so that filling the 1- to 3-electron ground states according to the Aufbau principle yields
+$\left\{\begin{split}\mathcal{E}_{N_{-}}(\Delta v^{\bm{\xi}}_{s})&=-\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2)^{2}}\\ \mathcal{E}_{N}(\Delta v^{\bm{\xi}}_{s})&=-2\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2)^{2}}\\ \mathcal{E}_{N_{+}}(\Delta v^{\bm{\xi}}_{s})&=-\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2)^{2}}\end{split}\right$.
+(S20)
+Inserting these expressions in Eq.
+S18 and using $\Delta n^{\bm{\xi}}_{s,N\pm p}=N\pm p-2\,n^{\bm{\xi}}_{s,N\pm p}$ where $n^{\bm{\xi}}_{s,N\pm p}$ is KS the density on site 0, one gets individual densities
+$\left\{\begin{split}\;&n^{\bm{\xi}}_{s,N_{-}}(\Delta v^{\bm{\xi}}_{s})=\frac{1}{2}+\frac{1}{4}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}\\ &n^{\bm{\xi}}_{s,N}(\Delta v^{\bm{\xi}}_{s})=1+\frac{1}{2}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}\\ &n^{\bm{\xi}}_{s,N_{+}}(\Delta v^{\bm{\xi}}_{s})=\frac{3}{2}+\frac{1}{4}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}\end{split}\right$.
+(S21)
+These allows to express the KS Fukui functions as
+$\left\{\begin{split}\;&f^{\bm{\xi}}_{s,-}(\Delta v^{\bm{\xi}}_{s})=\frac{1}{2}+\frac{1}{4}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}\\ &f^{\bm{\xi}}_{s,+}(\Delta v^{\bm{\xi}}_{s})=\frac{1}{2}-\frac{1}{4}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}\end{split}\right$.
+(S22)
+To obtain the ensemble KS response function, we first express ensemble density (on site 0) $n^{\bm{\xi}}$ as a function of the KS potential. Inserting Eq. (S21) into the definition
+$n^{\bm{\xi}}=\xi_{+}\,n^{\bm{\xi}}_{s,N_{+}}+\xi_{-}\,n^{\bm{\xi}}_{s,N_{-}}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)n^{\bm{\xi}}_{s,N}$ (S23)
+we obtain
+$n^{\bm{\xi}}(\Delta v^{\bm{\xi}}_{s})=1+\frac{1-\xi_{+}}{2}\frac{\Delta v^{\bm{\xi}}_{s}}{\sqrt{t^{2}+(\Delta v^{\bm{\xi}}_{s}/2})^{2}}$ (S24)
+Differentiation with respect to the KS potential yields the KS response function,
+$\chi_{s}^{\bm{\xi}}(\Delta v^{\bm{\xi}}_{s})=\frac{\partial n^{\bm{\xi}}(\Delta v_{s})}{\partial\Delta v_{s}}\Big|_{\Delta v_{s}=\Delta v^{\bm{\xi}}_{s}}=\frac{(1-\xi_{+})t^{2}}{2\left(t^{2}+(\Delta v^{\bm{\xi}}_{s}/2)^{2}\right)^{3/2}}$ (S25)
+Ensemble Hxc and KS potentials from the density – Inverting Eq. (S24) yields the ensemble KS potential as a function of ensemble density
+$\Delta v^{\bm{\xi}}_{s}(n^{\bm{\xi}})=\frac{2\,t\,(n^{\bm{\xi}}-1)}{\sqrt{(1-\xi_{+})^{2}-(n^{\bm{\xi}}-1)^{2}}}$ (S26)
+while the ensemble Hxc potential is defined as
+$v^{\bm{\xi}}_{\rm Hxc}[n]:=v^{\bm{\xi}}_{s}[n]-v^{\bm{\xi}}[n]$ (S27)
+where $v^{\bm{\xi}}[n]$ is the external potential reproducing the density $n$ for a given set of weights $\bm{\xi}$. For the Hubbard Dimer, the SOFT analog is
+$\Delta v^{\bm{\xi}}_{\rm Hxc}(n)=\Delta v^{\bm{\xi}}_{s}(n)-\Delta v^{\bm{\xi}}(n)$ (S28)
+with $\Delta v^{\bm{\xi}}(n)$ being the maximizing potential in the ensemble analog of Eq. (S9):
+$F^{\bm{\xi}}(n)=\underset{\Delta v}{\rm sup}\Big\{\xi_{-}\,E_{N_{-}}(\Delta v)+\xi_{+}\,E_{N_{+}}(\Delta v)+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\,E_{N}(\Delta v)+\Delta v(n-1)\Big\}$ (S29)
+No analytical expression is known from the above equation, but solving the optimization problem for given $n$ and $\bm{\xi}$ is straightforward, and can be done to arbitrary accuracy.
+Ensemble response and Hxc kernel – The ensemble response is defined as
+$\chi^{\bm{\xi}}[n^{\bm{\xi}}]({\bf r}):=\frac{\delta n^{\bm{\xi}}[v]}{\delta v({\bf r})}\Big|_{v=v^{\bm{\xi}}}$ (S30)
+with the Hubbard Dimer SOFT analog being
+$\chi^{\bm{\xi}}(n^{\bm{\xi}})=\frac{\partial n^{\bm{\xi}}(\Delta v)}{\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}$ (S31)
+Inserting ensemble density Eq. (S15) (see Eq. (S13)), we get
+$\chi^{\bm{\xi}}(n^{\bm{\xi}})=\frac{(\xi_{+}+\xi_{-})}{4}\frac{t^{2}}{\big[t^{2}+(\Delta v/2)^{2}\big]^{3/2}}-\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)E_{N}^{\prime\prime}$ (S32)
+where
+$E^{\prime\prime}_{N}=-\frac{4t}{3}\bigg(\big(w^{\prime\prime}-w\,\theta^{\prime 2}\big)\sin\left(\theta+\frac{\pi}{6}\right)+\big(w\,\theta^{\prime\prime}+\frac{3\nu}{2wt}\theta^{\prime}\big)\cos\left(\theta+\frac{\pi}{6}\right)\bigg)$ (S33)
+and
+$\left\{\begin{split}\;&w^{\prime\prime}=\frac{3}{4t}\frac{3+u^{2}}{w^{3}t}\\ &\theta^{\prime}=-\frac{3\nu}{w^{2}t\sin(3\theta)}\left(\frac{u}{w}-\frac{1}{2}\cos(3\theta)\right)\\ &\theta^{\prime\prime}=\frac{3}{2}\frac{1}{w^{2}t^{2}\sin(3\theta)}\left[\frac{1}{2}\cos(3\theta)-\frac{u}{w}+\frac{3}{2}\frac{\nu^{2}}{w^{2}}\left(12\frac{u}{w}-5\cos(3\theta)\right)\right]-3\theta^{\prime 2}\text{cot}(3\theta)\end{split}\right$.
+(S34)
+Finally, we can compute the Hxc kernel from ensemble Dyson equation
+$f_{\rm Hxc}^{\bm{\xi}}(n^{\bm{\xi}})=\Big(\chi^{\bm{\xi}}_{s}(n^{\bm{\xi}})\Big)^{-1}-\Big(\chi^{\bm{\xi}}(n^{\bm{\xi}})\Big)^{-1}$ (S35)
+using Eqs. (S25), (S26) and (S32).
+S-I.2 Taylor expansion in the low-correlation limit
+In all generality, the ensemble Hxc energy is related to the universal functional and the KS kinetic energy $T_{s}^{\bm{\xi}}[n]$ by
+$E^{\bm{\xi}}_{\rm Hxc}[n]:=F^{\bm{\xi}}[n]-T_{s}^{\bm{\xi}}[n]$ (S36)
+with the SOFT equivalent obtained by replacing functional dependence on a $\bf r$-dependent density by dependence on site occupation.
+In the low-correlation regime, one can expect a Taylor expansion in $U$ for the universal ensemble functional to be a good approximation:
+$\begin{split}F^{\bm{\xi}}(n)&:=\underset{\Delta v}{\rm sup}\Big\{\xi_{+}\,E_{N_{+}}(\Delta v)+\xi_{-}\,E_{N_{-}}(\Delta v)+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)E_{N}(\Delta v)+\Delta v(n-1)\Big\}\\ &=\underset{\Delta v}{\rm sup}\Big\{\xi_{+}U-(\xi_{+}+\xi_{-})\sqrt{t^{2}+(\Delta v/2)^{2}}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)E_{N}(\Delta v)+\Delta v(n-1)\Big\}\\ &=\sum_{i=0}^{\infty}\frac{U^{i}}{i!}\frac{\partial^{i}F^{\bm{\xi}}(n)}{\partial U^{i}}\Big|_{U=0}\end{split}$ (S37)
+As the true system becomes identical to its KS analog in the 0-correlation limit,
+$F^{\bm{\xi}}(n)\underset{U\rightarrow 0}{=}T^{\bm{\xi}}_{s}(n)$ (S38)
+we can express the Hxc energy from the Taylor expansion of $F^{\bm{\xi}}(n)$:
+$E^{\bm{\xi}}_{\rm Hxc}(n)=\sum_{i=1}^{\infty}\frac{U^{i}}{i!}\frac{\partial^{i}F^{\bm{\xi}}(n)}{\partial U^{i}}\Big|_{U=0}$ (S39)
+The first term yields the ensemble Hartree-exchange energy, which we express from Eq. (S37) as
+$E^{\bm{\xi}}_{\rm Hx}(n)=U\frac{\partial F^{\bm{\xi}}(n)}{\partial U}\Big|_{U=0}=U\left[\xi_{+}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\frac{\partial E_{N}(\Delta v)}{\partial U}\right]_{\Delta v=\Delta v^{\bm{\xi}}(n,U=0)}$ (S40)
+where we explicitly indicated the $U$-dependence of $\Delta v^{\bm{\xi}}(n)$, that is the external potential reproducing the trial ensemble density $n$ for a given set of weights.
+Differentiating the polynomial Eq.
+(S3) with respect to $U$ yields the expression
+$\frac{\partial E_{N}}{\partial U}\Big|_{U=0}=\frac{\nu^{2}+1/2}{1+\nu^{2}}$ (S41)
+From the identity of KS and true systems in the 0-correlation limit,
+$\nu=\frac{\Delta v^{\bm{\xi}}}{2t}\underset{U\rightarrow 0}{=}\frac{\Delta v^{\bm{\xi}}_{s}}{2t}$ (S42)
+we can insert KS potential Eq. (S26) and get
+$E^{\bm{\xi}}_{\rm Hx}(n)=\frac{U}{2}\left[1+\frac{\xi_{+}-\xi_{-}}{2}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\frac{(n-1)^{2}}{(1-\xi_{+})^{2}}\right]$ (S43)
+Using this term alone corresponds to the EEXX approximation in main text. The second order term in the Taylor expansion, which constitutes the PT2 approximation to the correlation energy, reads
+$\begin{split}\frac{\partial^{2}F^{\bm{\xi}}(n)}{\partial U^{2}}\Big|_{U=0}=\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\bigg[&\frac{\partial^{2}E_{N}(\Delta v)}{\partial U^{2}}\Big|_{\Delta v=\Delta v^{\bm{\xi}}(U=0)}\\ &+\frac{\partial^{2}E_{N}(\Delta v)}{\partial U\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}(U=0)}\frac{\partial\Delta v^{\bm{\xi}}(U)}{\partial U}\Big|_{U=0}\bigg]\end{split}$ (S44)
+Differentiating the polynomial Eq. (S3) appropriately yields
+$\frac{\partial^{2}E_{N}(\Delta v)}{\partial U^{2}}\Big|_{\Delta v=\Delta v^{\bm{\xi}}(U=0)}=-\frac{1}{2t}\frac{\nu^{2}+1/4}{(1+\nu^{2})^{5/2}}$ (S45)
+and
+$\frac{\partial^{2}E_{N}(\Delta v)}{\partial U\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}(U=0)}=\frac{\nu}{2t(1+\nu^{2})^{2}}$ (S46)
+To obtain the $U$-derivative of the ensemble potential, note the Lieb maximization principle Eq. (S29) implies
+$\frac{\partial F^{\bm{\xi}}(n,\Delta v)}{\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}=0$ (S47)
+Inserting Eq. (S37) yields
+$(\xi_{+}+\xi_{-})\frac{\partial\epsilon_{H}(\Delta v)}{\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\frac{\partial E_{N}(\Delta v)}{\partial\Delta v}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}+n-1=0$ (S48)
+and differentiating again with respect to $U$:
+$\begin{split}&(\xi_{+}+\xi_{-})\frac{\partial^{2}\epsilon_{H}(\Delta v)}{\partial\Delta v\partial U}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\frac{\partial^{2}E_{N}(\Delta v)}{\partial\Delta v\partial U}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}\\ &+\left[(\xi_{+}+\xi_{-})\frac{\partial^{2}\epsilon_{H}(\Delta v)}{\partial\Delta v^{2}}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\frac{\partial^{2}E_{N}(\Delta v)}{\partial\Delta v^{2}}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}\right]\frac{\partial\Delta v^{\bm{\xi}}(n)}{\partial U}=0\end{split}$ (S49)
+Using Eqs.
+(S19,S45,S46) and differentiating polynomial Eq.
+(S3) with respect to $\Delta v$ twice to compute $\frac{\partial^{2}E_{N}(\Delta v)}{\partial\Delta v^{2}}\Big|_{\Delta v=\Delta v^{\bm{\xi}}}$ we finally obtain
+$\frac{\partial\Delta v^{\bm{\xi}}(n)}{\partial U}\Big|_{U=0}=\frac{\nu}{\sqrt{1+\nu^{2}}}\frac{2-(\xi_{-}+3\xi_{+})}{2(1-\xi_{+})}$ (S50)
+We can now insert Eqs. (S45,S46,S50) in Eq. (S44) to get the second order term in the Taylor expansion of the universal functional
+$\frac{\partial^{2}F^{\bm{\xi}}(n)}{\partial U^{2}}\Big|_{U=0}=\frac{2-\xi_{-}-3\xi_{+}}{16t}\left(\frac{(n-1)^{2}}{(1-\xi_{+})^{2}}\frac{1-2\xi_{-}-3\xi_{+}}{1-\xi_{+}}-1\right)\left[\frac{(1-\xi_{+})^{2}-(n-1)^{2}}{(1-\xi_{+})^{2}}\right]^{3/2}$ (S51)
+which completes the PT2 approximation to the ensemble Hxc energy.
+$E^{\bm{\xi}}_{\rm Hxc}(n)\underset{PT2}{\approx}U\frac{\partial F^{\bm{\xi}}(n)}{\partial U}\Big|_{U=0}+\frac{U^{2}}{2}\frac{\partial^{2}F^{\bm{\xi}}(n)}{\partial U^{2}}\Big|_{U=0}$ (S52)
+S-I.3 High density limit and the PADE interpolant
+To derive the high correlation limit corresponding to $u=U/2t\rightarrow\infty$, we reexpress the polynomial Eq. (S3) with $\tilde{E}_{N,i}=E_{N,i}/U$, $\Delta\tilde{v}=\Delta v/U$, and $u$:
+$\tilde{E}_{N,i}^{3}=-\frac{1}{u^{2}}+\left(\frac{1}{u^{2}}-1+\Delta\tilde{v}^{2}\right)\tilde{E}_{N,i}+2\tilde{E}_{N,i}^{2}$ (S53)
+in the high-$u$ limit, it becomes
+$\tilde{E}_{N,i}^{3}\underset{u\rightarrow\infty}{=}\left(\Delta\tilde{v}^{2}-1\right)\tilde{E}_{N,i}+2\tilde{E}_{N,i}^{2}$ (S54)
+which can straightforwardly be put under the form of a second order polynomial, with lowest solution (i.e. the ground-state) given by
+$\tilde{E}_{N}\underset{u\rightarrow\infty}{=}{\rm min}\big\{0,1-\Delta\tilde{v}\big\}$ (S55)
+This, together with the knowledge of $u\rightarrow\infty$ limit for the non-interacting one-electron energy,
+$\frac{\epsilon_{H}(\Delta\tilde{v})}{U}\underset{u\rightarrow\infty}{=}-\left|\frac{\Delta\tilde{v}}{2}\right|$ (S56)
+allows to rexpress the ensemble functional of Eq. (S37)
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty}{=}\underset{\Delta\tilde{v}}{{\rm sup}}\left\{\xi_{+}-|\Delta\tilde{v}|\frac{\xi_{+}+\xi_{-}}{2}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right){\rm min}\big\{0,1-\Delta\tilde{v}\big\}+\Delta v(n-1)\right\}$ (S57)
+To make further progress, one has to consider several $\Delta\tilde{v}$ ranges separately. For $|\Delta\tilde{v}|\leq 1$,
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty,|\Delta\tilde{v}|\leq 1}{=}\underset{|\Delta\tilde{v}|\leq 1}{{\rm sup}}\left\{\xi_{+}-|\Delta\tilde{v}|\frac{\xi_{+}+\xi_{-}}{2}+\Delta v(n-1)\right\}$ (S58)
+which is further decomposed into
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty,|\Delta\tilde{v}|\leq 1}{=}\left\{\begin{split}\quad&{\rm sup}\left\{\xi_{+}\,,\,\frac{\xi_{+}-\xi_{-}}{2}-(n-1)\right\}\quad{\rm if}\;-1\leq|\Delta\tilde{v}|\leq 0\\ &{\rm sup}\left\{\xi_{+}\,,\,\frac{\xi_{+}-\xi_{-}}{2}+(n-1)\right\}\quad{\rm if}\;0\geq|\Delta\tilde{v}|\geq 1\end{split}\right$.
+(S59)
+so that we can finally state
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty,|\Delta\tilde{v}|\leq 1}{=}{\rm sup}\left\{\xi_{+}\,,\,\frac{\xi_{+}-\xi_{-}}{2}+|n-1|\right\}$ (S60)
+Shifting to $|\Delta\tilde{v}|\geq 1$,
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty,|\Delta\tilde{v}|\geq 1}{=}\underset{|\Delta\tilde{v}|\geq 1}{{\rm sup}}\left\{\xi_{+}-|\Delta\tilde{v}|\frac{\xi_{+}+\xi_{-}}{2}+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\big(1-\Delta\tilde{v}\big)+\Delta v(n-1)\right\}$ (S61)
+we first consider negative values
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty}{=}\underset{\Delta\tilde{v}\leq-1}{{\rm sup}}\Big\{\Delta\tilde{v}\big(n-\xi_{+}\big)\Big\}+1-\frac{\xi_{+}+\xi_{-}}{2}\quad{\rm if}\;\Delta\tilde{v}\leq-1$ (S62)
+To make progress, we take insight from non-interacting $v$-representability: the analytical expression of $T^{\bm{\xi}}_{s}(n)$ reads
+$\begin{split}T^{\bm{\xi}}_{s}(n)&=\underset{\Delta v}{\rm sup}\Big\{\xi_{-}\,\mathcal{E}_{N_{-}}(\Delta v)+\xi_{+}\,\mathcal{E}_{N_{+}}(\Delta v)+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\,\mathcal{E}_{N}(\Delta v)+\Delta v(n-1)\Big\}\\ &=\underset{\Delta v}{\rm sup}\Big\{(\xi_{-}+\xi_{+})\,\epsilon_{H}(\Delta v)+\xi_{+}\,U+\left(1-\frac{3\xi_{+}+\xi_{-}}{2}\right)\,\epsilon_{H}(\Delta v)+\Delta v(n-1)\Big\}\\ &=\underset{\Delta v}{\rm sup}\Big\{-2\sqrt{t^{2}+(\Delta v/2)^{2}}\,(1-\xi_{+})+\Delta v(n-1)\Big\}\end{split}$ (S63)
+inserting the maximizing potential expression, that is KS potential Eq. (S26), one finally gets
+$T^{\bm{\xi}}_{s}(n)=-2t\sqrt{(1-\xi_{+})^{2}-(n-1)^{2}}$ (S64)
+which is well-defined only when
+$1-\xi_{+}\geq n-1\geq\xi_{+}-1$ (S65)
+Thus, $n-\xi_{+}\geq 0$ and
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty}{=}\frac{\xi_{+}-\xi_{-}}{2}-(n-1)\quad{\rm if}\;\Delta\tilde{v}\leq-1$ (S66)
+Moving to $\Delta\tilde{v}\geq 1$, we have
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty}{=}\underset{\Delta\tilde{v}\leq-1}{{\rm sup}}\Big\{\Delta\tilde{v}\big(n-2+\xi_{+}\big)\Big\}+1-\frac{\xi_{+}+\xi_{-}}{2}\quad{\rm if}\;\Delta\tilde{v}\geq 1$ (S67)
+As from Eq. (S65), we have $n-2+\xi_{+}\leq 0$, we obtain
+$\frac{F^{\bm{\xi}}(n)}{U}\underset{u\rightarrow\infty}{=}\frac{\xi_{+}-\xi_{-}}{2}+(n-1)\quad{\rm if}\;\Delta\tilde{v}\geq 1$ (S68)
+which can be rewritten
+$F^{\bm{\xi}}(n)\underset{u\rightarrow\infty}{=}U\;{\rm sup}\left\{\xi_{+}\,,\,\frac{\xi_{+}-\xi_{-}}{2}+|n-1|\right\}$ (S69)
+Knowing the high-$U$ limit and the low-$U$ second-order Taylor expansion, we can build the following Padé approximant to interpolate between them:
+$E_{\rm Hxc}^{\bm{\xi}}(n)\approx a^{\bm{\xi}}(n)U+\frac{b^{\bm{\xi}}(n)U^{2}}{1+c^{\bm{\xi}}(n)U}$ (S70)
+where all terms are given by
+$\left\{\begin{split}\;a^{\bm{\xi}}(n)&=\frac{\partial F^{\bm{\xi}}(n)}{\partial U}\Big|_{U=0}\\ b^{\bm{\xi}}(n)&=\frac{1}{2}\frac{\partial^{2}F^{\bm{\xi}}(n)}{\partial U^{2}}\Big|_{U=0}\\ c^{\bm{\xi}}(n)&=\frac{b^{\bm{\xi}}(n)}{\gamma^{\bm{\xi}}(n)-a^{\bm{\xi}}(n)}\\ \gamma^{\bm{\xi}}(n)&={\rm sup}\left\{\xi_{+}\,,\,\frac{\xi_{+}-\xi_{-}}{2}+|n-1|\right\}\end{split}\right$.
+(S71)
+Appendix S-II Numerical details and complementary results
+S-II.1 Computational details
+All calculations are done at the exact ensemble density $n^{\bm{\xi}}$ determined from the system parameters using Eqs.
+(S13) to (S15).
+See definitions in and below Eq.
+(S5).
+Of course, calculations using scaling functions are performed in the zero-weight limit where the ensemble density reduces to the 2-electron ground state density given in Eq.
+(S13).
+Various values of $U$ and $\Delta v$ will be considered, but in all cases we set $t=1$.
+Calculations using full-ensemble functionals – For arbitrary weights $(\xi_{-},\xi_{+})$, Fukui functions are computed from the working equation:
+$\begin{split}f_{\pm}&=\big(1+\chi^{\bm{\xi}}f_{\rm Hxc}^{\bm{\xi}}\big)f^{\bm{\xi}}_{s,\pm}-\chi^{\bm{\xi}}f_{\rm Hxc}^{\bm{\xi}}\frac{n^{\bm{\xi}}}{2}\\ &\quad\pm\;\chi^{\bm{\xi}}\sum_{\lambda=+,-}\Big(\delta_{\lambda\pm}\mp\frac{\xi_{\lambda}}{2}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}(n)}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}.\end{split}$ (S72)
+with the shorthand notation $f^{\bm{\xi}}:=f^{\bm{\xi}}(n^{\bm{\xi}})$ for any function of the density (one site 0).
+The KS Fukui functions $f^{\bm{\xi}}_{s,\pm}$ are computed analytically from the ensemble density using Eq.
+(S22).
+The ensemble response $\chi^{\bm{\xi}}$ is computed from ensemble Dyson Eq.
+(S35) using the analytical KS response defined through Eq.
+(S25) and the Hxc kernel $f^{\bm{\xi}}_{\rm Hxc}$.
+The latter is evaluated through analytical differentiation of the relevant density functional approximation of Hxc energy.
+Weight-derivatives of the Hxc potential are obtained in the same manner.
+Note however that our SOFT convention to express everything as a function of density on site 0 brings some sign changes compared to general EDFT expressions: from the first line of Eq.
+(S37) and its non-interacting analog, it is clear our external and KS potentials have a flipped sign compared to general EDFT expressions using functionals
+$\left\{\begin{split}\Delta v^{\bm{\xi}}_{s}(n)&=\frac{\partial T^{\bm{\xi}}_{s}(n)}{\partial n}\\ \Delta v^{\bm{\xi}}(n)&=\frac{\partial F^{\bm{\xi}}(n)}{\partial n}\end{split}\right$. (S73)
+From the Hxc energy and potential definitions through Eq.
+(S36), which carries straightforwardly to SOFT, and through Eq.(S28) respectively, we see the sign flip carries to the relation between ensemble Hxc energy and potential:
+$\begin{split}\Delta v^{\bm{\xi}}_{\rm Hxc}(n)&=\Delta v^{\bm{\xi}}_{s}(n)-\Delta v^{\bm{\xi}}(n)\\ &=\frac{\partial}{\partial n}\big[T^{\bm{\xi}}_{s}(n)-F^{\bm{\xi}}(n)\big]\\ &=-\frac{\partial E^{\bm{\xi}}_{\rm Hxc}(n)}{\partial n}\end{split}$ (S74)
+Regarding the kernel, owing to the KS and physical response function definitions Eqs.
+(S25,S31), differentiating the above expression for the Hxc potential with respect to $n$ and using ensemble Dyson Eq.
+(S35) gives
+$f^{\bm{\xi}}_{\rm Hxc}(n)=\frac{\partial\Delta v^{\bm{\xi}}_{\rm Hxc}(n)}{\partial n}=-\frac{\partial^{2}E^{\bm{\xi}}_{\rm Hxc}(n)}{\partial n^{2}}$ (S75)
+Calculations using scaling functions – The exact ground-state functional is always used in combination with scaling functions, everything being evaluated at the exact ground-state density.
+All scaling functions are evaluated from relevant Hxc potential approximations at the given weights $\bm{\xi}$ and for 0-weights.
+$s(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}{v_{\rm Hx}[n^{\bm{\xi}}]}$ (S76)
+We stress again potentials are to be evaluated at the relevant ensemble density $n^{\bm{\xi}}$ to build the scaling function.
+The purely Hx scaling (or EEXX scaling) is computed from the ratio of Hx potentials (obtained by inserting Eq.
+(S43) in Eq.
+(S74)).
+At the EEXX level for the scaling function, the density dependences simplify out.
+For the mono-scaling at PT2 level, the Hxc potential is obtained by differentiation of Eq.
+(S52) with respect to $n$ (see again Eq.
+(S74)).
+For the double-scaling at PT2 level, the Hx scaling function is given by EEXX potential, while the correlation scaling $s_{c}(\bm{\xi})$ is built from the ratio of correlation potentials obtained by differentiating the second term of Eq.
+(S52) with respect to $n$ (see again Eq.
+(S74)).
+S-II.2 EEXX and PT2 in the low correlation regime
+When correlation is low, one expects EEXX to capture most of the physics.
+Its performance in predicting Fukui functions in this regime is shown in Fig.
+S1 for $U=0.75$ and 0 weights.
+It is in good agreement with exact results and the neglect of weight-derivative terms in Eq.
+(S72) is shown for comparison.
+The latter result uses the exact ground-state functional to evaluate exactly all other terms.
+Figure S1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset).
 Full and dashed lines are $f_{+}$ and $f_{-}$ calculations respectively.
 Exact results (yellow lines) are compared to EEXX (blue lines) and to the neglect of weight-derivative contributions (grey lines).
-In the weak correlation regime $U/t=0.75$, the EEXX functional closely reproduces exact results as seen on Fig.
-1, with the effect of neglecting Hxc weight derivatives shown in grey.
-The improvement is clear, making this first application of EEXX to compute densities promising.
-These encouraging results suggest using the scaling function
-$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hxc}[n]\,,\qquad s_{\rm Hx}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}{v_{\rm Hx}[n^{\bm{\xi}}]}$ (17)
-Note the ensemble density is fed to the ground-state potential.
-It is important to define the scaling in this way to ensure evaluating the weight-derivative of the scaling function (in the 0-weight limit) does not require knowing the weight-derivative of the ensemble density.
-While using this scaling function misses correlation at the ensemble level, note some measure of it is still brought into the approximate potential through the ground-state Hxc potential (see Eq.
-(17)).
-For $U/t=0.75$, there is negligible difference between the EEXX scaling method (not shown) and full EEXX DFA, which could be expected as correlation barely contributes in this regime.
-Figure 2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset).
-We compare EEXX ensemble DFA (blue) to EEXX scaling (red) and neglect of weight derivative terms (grey).
-Exact results are shown in orange.
-Intermediate correlation regime – We expect this to change by turning to higher (but moderate) correlation $U/t=1.5$, results being shown on Fig.
-2.
-For external potentials below $\Delta v=2U/t$, where the correlation energy becomes important, we see EEXX deviates from the exact result.
-It still improves substantially over the neglect of weight-derivative terms for $f_{-}$, but not for $f_{+}$: the weight-derivative terms only bring a small correction here, which is overestimated in EEXX.
-Turning to the scaling strategy, we ask: Is ground-state correlation a reasonable approximation of its ensemble analog?
-Results show dressing the ground-state correlation potential with a purely Hx ensemble scaling brings a modest improvement over full EEXX for $f_{-}$, but prediction of $f_{+}$ actually worsens.
-These results show it is necessary to go beyond EEXX when correlation takes over.
-Moreover, we have to understand if the failure of Hx scaling function comes from an inherent discrepancy between ground-state and ensemble correlation, making the weight-dependent dressing strategy impractical, or if it can be fixed.
-To deal with situations where EEXX fails, second order (Görling-LeviGörling and Levy (1994)) perturbation theory has been applied to ensemble DFT, yielding an orbital-dependent DFA beyond EEXXYang (2021).
-We transpose this in the Hubbard dimer through second-order Taylor expansion in $U$:
-$E^{\bm{\xi}}_{\rm c}[n]\overset{\rm PT2}{\approx}\frac{U^{2}}{2}\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}$ (18)
-Compared to EEXX-only on Fig. 3, it shows much better agreement with exact Fukui functions.
-Figure 3: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset).
+The situation changes when increasing correlation up to $U=1.5$ (keeping 0 weights) as shown on Fig.
+S2.
+A noticeable discrepancy between EEXX and exact results appears for low external potentials $\Delta v<2U$, where the correlation energy is the highest.
+Note however the improvement over neglect of weight-derivative terms is still noticeable for $f_{-}$.
+Inclusion of correlation in the PT2 DFA allows to recover a close agreement with exact results in all the $\Delta v$ range.
+Figure S2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset).
 We compare EEXX ensemble DFA (blue) to PT2 (green) and neglect of weight derivative terms (grey).
 Exact results are shown in orange.
-Moreover, comparison of functional forms of EEXX and PT2 shows where our Hx-only scaling is failing: The ratio of PT2 ensemble and ground-state correlation potentials is an explicit function of the weights and of the ensemble density:
-$s_{\rm c,PT2}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm c,PT2}[n^{\bm{\xi}}]}{v_{\rm c,PT2}[n^{\bm{\xi}}]}=s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})$ (19)
-while in the Hx ratio Eq.
-(17) all dependence on the ensemble density simplifies out, leaving only explicit weight dependences.
-This disparity in complexity for the Hx and correlation part suggests using a different scaling function for each, which reads at the PT2 level:
-$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})\;v_{\rm c}[n]$ (20)
-with exact ensemble density dependence shown explicitly.
-Note if the correlation term was neglected all-together, leaving only the first term, the potential would be equal to that of EEXX (in the 0-weight limit).
-But at the PT2 level, it dresses the exact ground-state correlation with an ensemble-density dependence.
-The question now becomes: Does this double scaling at the PT2 level improve over the full PT2 EDFA?
-On Fig.
-4, we compare their performance in computing $f_{+}$ for $U/t$ up to 2.5.
-We can see that the PT2 scaling accurately matches exact results even above $U=1.5$, when both EEXX and PT2 fail.
-This shows the scaling function strategy holds promise, but it is crucial to treat the Hx and correlation parts separately.
-Figure 4: Fukui function $f_{+}$ in the low correlation regime for null ensemble weights (parameters in the red inset).
-For increasing $U$, we compare our EEXX EDFA (blue) to PT2 EDFA (green) and the PT2 scaling strategy (red).
-Exact results (orange) and neglect of weight-derivatives (grey) are plotted for reference.
-Strong correlation – Learning from high- and low-correlation limits of the EDFT proved to be an outstanding approach for excitation energies in difficult situationsGould et al.
-(2023), with a versatile EDFA obtained from interpolating between the two regimes.
-We put it to the test for Fukui functions by building a Padé interpolation between the PT2 (low $U$) and $U/t\rightarrow\infty$ limits of the ensemble Hxc functional,
-$E_{\rm Hxc}^{\bm{\xi}}[n]=a^{\bm{\xi}}[n]\,U+\frac{b^{\bm{\xi}}[n]\,U^{2}}{1+c^{\bm{\xi}}[n]\,U}$ (21)
-where
-$\begin{split}a^{\bm{\xi}}[n]&=\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}\\ b^{\bm{\xi}}[n]&=\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}\\ c^{\bm{\xi}}[n]&=\frac{b^{\bm{\xi}}[n]}{\gamma^{\bm{\xi}}[n]-a^{\bm{\xi}}[n]}\end{split}$ (22)
-and the strong-correlation-dominant term
-$\gamma^{\bm{\xi}}[n]=\lim_{U/t\rightarrow\infty}E_{\rm Hxc}^{\bm{\xi}}[n]={\rm sup}\Big\{\xi_{+},\frac{\xi_{+}-\xi_{-}}{2}+\big|n-1\big|\Big\}$ (23)
-Figure 5: Fukui functions in strong correlation regime for non-zero weights (parameters in the red inset).
-We compare the Padé EDFA (green) to EEXX prediction (blue).
-Exact results (orange) and neglect of the weight-derivative terms (grey) are plotted for reference.
-This EDFA performs poorly (not shown) as its discontinuous behavior (through the sup function) and density-derivative discontinuity in |n1||n-1| means it lacks cross-derivatives (in $n$ and $\xi_{\pm}$) contributions to $\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n^{\bm{\xi}}}$ that are important even in the high correlation regime.
-Even if Eq.
-(23) is formally exact in the strictly-correlated regime ($U\rightarrow\infty$), it is not representative of the strongly-correlated regime (typically $U\in[5,10]$).
-To remedy this, the sharp sup function and absolute value were each replaced by a smoothing function whose stiffness parameters increase with $U$, thus yielding cross derivative terms while still being compatible with the exact $U/t\rightarrow\infty$ limit.
-To set the stiffness parameters in a way consistent with what could be done for a real system, we only considered the symmetric case of 2-electrons Hubbard dimer, i.e.
-$\Delta v=0$ and $\bm{\xi}=0$, which is the analog of the uniform electron gas in this context.
-The stiffness parameters were set so that our Padé approximation reproduces best the (ground-state) Hxc kernel.
-The obtained EDFA is in principle applicable for any choice of weights, even if it was smoothed by learning from the ground-state symmetric case.
-It is thus most interesting to test its performance for non-zero weights.
-[…]
-This trend carries over to higher correlation, as shown on Fig.
-6 for $U/t=10$ and $\xi=0.2$, which is remarkable as the stifness parameters smoothing the $U\rightarrow\infty$ discontinuities in $\gamma^{\bm{\xi}}[n]$ where fixed using only the ground-state, symmetric limit of the Hxc functional.
-Figure 6: Fukui functions in very strong correlation regime for non-zero ensemble weights (parameters in the red inset). See Fig. 5 for labels.
-Summary and outlook– After introducing the (in principle) exact expression of Fukui functions within Nc EDFT, we discussed the equivalence between weight-derivatives of the Hxc potential and derivative discontinuities in the Hxc kernel.
-The exact relations introduced by the Authors between individual true physical and non-interacting KS densitiesFromager (2025) and responsesDupuy and Fromager (2025) are thus a promising framework to access properties beyond excitation energies within EDFT.
-By considering the conventional $N$-electron ground-state DFT limit of the theory, we showed how dressing the regular (ground-state) Hxc functional with a weight-dependent scaling function allows to recycle practical (semi-)local DFAs while accurately accounting for the ensemble Hxc potential weight-derivatives term, as illustrated on Fukui functions of the Hubbard Dimer.
-It was instrumental to the performance of this strategy that separate scaling functions would be applied to the Hx and correlation parts of the ground-state potential, as their ensemble analogs had starkly different dependence on ensemble weights and density.
-This is in line with analysis of low- and high-correlation limits of EDFTGould et al.
-(2023), predicting a dominant dependence on the ensemble density rather than on the weights in the latter case.
-When used properly, the scaling function approach extended the domain of applicability of its parent approximation by exploiting an accurate ground-state functional.
-We also showed building a full ensemble DFA by interpolating between known low- and high-correlation limits performed very well, provided one paid attention to avoid discontinuous behavior in the functional.
-Smoothing over these discontinuities by learning from the symmetric ground state proved very effective, echoing recent work on the uniform electron gas as a basis to develop excited state density functionalsLoos (2025).
-These findings will foster DFA development in realistic systems.
-As additional strategies, one could also include exact constraints satisfied by ensembles in the design of scaling functions, such as the invariance of individual-state physical densities with respect to weights (see Eq.
-(8)), which is currently explored.
-As our formalism applies to any (neutral or charged) excitation, it can bring insights into processes that PPLB could not grasp, notably excited state reactivity through the relevant Fukui functionsAn et al.
-(2024).
-This is left for future work.
-Acknowledgements - This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).
+S-II.3 Scaling functions
+Inadequacy of an Hxc energy-based scaling function – Using the ratio of Hxc ensemble and ground-state energies at the ensemble density is a poor strategy to predict the Fukui function.
+$s(\bm{\xi})=\frac{E^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]}{E_{\rm Hxc}[n^{\bm{\xi}}]}$ (S77)
+The reason is that they encode widely-different weight-dependence, as seen on Fig.
+S3 where we compare the two definitions for the scaling function at $U/t=2.5$.
+The weight $\xi_{-}$ is kept fixed at 0, while $\xi_{+}$ is varied.
+Different values of the external potential, producing different ensemble densities, are considered.
+In all cases, it is seen that the scaling functions using energies and potentials vary in an almost opposite way, making the former unsuited to encode weight-dependence of the Hxc potential.
+Figure S3: Profiles of the scaling function as a function of $\xi_{+}$ defined from the Hxc energy ratio (full lines) compared to the ratio of potentials (dashed lines).
+The weight $\xi_{-}$ is kept at 0.
+Three values of the external potential are considered.
+Improvement brought by double scaling at the PT2 level – To understand the impact of the double-scaling, we compare on Fig.
+S4 the ensemble correlation potential predicted by the full-PT2 DFA to the one obtained by scaling the ground-state correlation potential using to the ratio of PT2 ensemble and ground-state correlation potentials.
+Each of the four panels correspond to a different correlation strength $U/t$.
+On each panel, in the 0-weight case, the scaling strategy is exact by construction, while the full PT2 correlation potential exhibits increasing errors with $U/t$ when the system is strongly asymmetric.
+Results for $\xi=\xi_{+}=\xi_{-}=0.05$ show the scaled potential matches closely the exact ensemble one, even for large $U/t$, with the brunt of the error around the symmetric case.
+This shows the weight-derivatives of the Hxc potential will be modeled more accurately for large $\Delta v$ thanks to the scaling approach, while they will not improve nor deteriorate results around $\Delta v=0$.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure S4: Ensemble correlation energy as a function of ensemble density.
+Exact results (full lines) are compared to full PT2 potential (long-dashed lines) and to the PT2 double scaling function applied to the ground state potential (short-dashed lines).
+Two values are considered for the weights, taken to be equal $\xi=\xi_{+}=\xi_{-}$.
+Each panel corresponds to a specific value of $U$.
+This is illustrated on Fig. S5 for $U/t=1.5$, comparing exact weight-derivatives of the ensemble Hxc energy to the full PT2 EDFA and the PT2-derived double scaling.
+Figure S5: Weight-derivatives of the ensemble Hxc potential in the 0-weight limit.
+Exact results (full lines) are compared to full PT2 EDFA (long-dashed lines) and PT2 scaling of ground state DFA (short-dashed lines).
+S-II.4 Smoothing of Pade interpolant
+Although $U\,\gamma^{\bm{\xi}}(n)$ defined through Eq.
+(S71) is the exact limit of $E^{\bm{\xi}}_{\rm Hxc}(n)$ for $u\rightarrow\infty$, its discontinuous behavior is a poor approximation of the strong correlation regime (say, $U/t>5$.) On Figs.
+S6 to S9, the right panels show it does not reproduce Fukui functions accurately for $\Delta v<U$, considering to sets of weights and two values of $U$.
+The discontinuities in $\gamma^{\bm{\xi}}$ derivatives are reflected in the Fukui functions, at $\Delta v=0$ for the 0-weight case, and displaced around $2\xi\,U$ for $\xi=\xi_{+}=\xi_{-}=0.2$.
+The left panels show the Padé approximant error comes from a missing contribution to the weight-derivative of the Hxc potential for low $\Delta v$.
+Its prediction for the Hxc kernel, shown in middle panels, is also poor in this range.
+It is reasonable to assume that for high-but-finite $U$, the Hxc energy has continuous derivatives with respect to the density and weights, which yield cross-derivative contributions to $\frac{\partial v^{\bm{\xi}}_{\rm Hxc}(n)}{\partial\xi_{\pm}}|_{n=n^{\bm{\xi}}}$.
+Similarly, it would induce important variations of the kernel around the derivative discontinuities of the Padé approximant that we observe.
+We thus replace the original expression for $\gamma^{\bm{\xi}}(n)$ by a smoothed version:
+$\begin{split}\bar{\gamma}^{\bm{\xi}}(n)&=\eta(n-1)+\frac{1}{2}\left(\xi_{+}-\xi_{-}+\frac{1}{k_{\xi}}\ell n(1+e^{k_{\xi}[\xi_{+}+\xi_{-}-2*\eta(n-1)]}\right)\\ \eta(n-1)&=\frac{2}{k_{n}}\ell n\left(1+e^{k_{n}*(n-1)}\right)-(n-1)\end{split}$ (S78)
+where the parameters $k_{n}$ and $k_{\xi}$ control the stiffness of the function, and should be functions of $U$.
+This is an opportunity to explore how conditions on the functional in familiar regimes can guide DFA development.
+To do so, we set $k_{n}$ and $k_{\xi}$ so that $f^{\bm{\xi}}_{\rm Hxc}(n)$ is correctly reproduced in the symmetric case ($\Delta v=0$), corresponding to the uniform electron gas limit, for zero weights.
+This is illustrated on middle panels of Figs.
+S6 and S8.
+Multiple couple values $(k_{n},k_{\xi})$ can satisfy this requirement.
+As an additional constraint, we minimize the difference between the exact and Padé slopes of $\frac{\partial\Delta v^{\bm{\xi}}_{\rm Hxc}(n)}{\partial\xi_{\pm}}$ for $\Delta v=0$ and zero weights, as seen on left panels of aforementioned figures.
+This yields the parameters $(k_{n}=64,k_{\xi}=15)$ for $U=5$ and $(k_{n}=130,k_{\xi}=25)$ for $U=10$.
+The improvement over the Padé approximation with discontinuous derivatives can be seen on right panels of Figs.
+S6 and S8 for the prediction of Fukui functions.
+More interesting, in our opinion, is the impact outside the 0-weight limit.
+One can see the smoothed Padé approximant improves substantially the Hxc kernel, weight-derivatives of Hxc potential, and resulting Fukui functions for $\xi=\xi_{+}=\xi_{-}=0.2$ on Figs.S7 and S9.
+This is a promising result for transferability as only the 0-weight limit was used to improve the Padé approximant.
+Of course, other constraints could be considered, and our choice of smoothing functions is entirely arbitrary.
+In that regard, performing a Taylor expansion in $1/u$ around $1/u=0$, in the spirit of ISIGould et al.
+(2023) would be informative in realistic systems.
+The present constructions simply serves to demonstrate care must be taken with discontinuous derivatives in ensemble DFAs.
+Refer to caption Refer to caption Refer to caption
+Figure S6: Effect of smoothing the strictly-correlated limit in the Padé approximant. Left panel: Weight-derivatives of the Hxc potential. Middle panel: Hxc kernel. Right panel: Fukui functions.
+Refer to caption Refer to caption Refer to caption
+Figure S7: Same as Fig. S6 for non-zero weights $\xi=\xi_{+}=\xi_{-}$.
+Refer to caption Refer to caption Refer to caption
+Figure S8: Same as Fig. S6 with $U/t=10$.
+Refer to caption Refer to caption Refer to caption
+Figure S9: Same as Fig. S6 with $U/t=10$ and $\xi=\xi_{+}=\xi_{-}=0.2$.